# Procedure Learning via Regularized Gromov-Wasserstein Optimal Transport

### v1

@@ 5 Conclusion
We present a self-supervised procedure learning framework for recognizing key steps and their order from a collection of unlabeled procedural videos. In particular, our approach leverages a fused Gromov-Wasserstein optimal transport formulation with a structural prior to learn frame-to-frame correspondences between videos while tackling order variations, background/redundant frames, and repeated actions. Furthermore, we empirically examine possible trivial solutions when optimizing solely for the above temporal alignment term and integrate a contrastive regularization term which encourages various frames to be mapped to various points in the embedding space to prohibit the collapse to degenerate solutions. Lastly, extensive evaluations on large-scale egocentric (i.e., EgoProceL) and third-person (i.e., ProCeL and CrossTask) benchmarks are performed to show our superior performance over previous methods, including OPEL which is based on a classical Kantorovich optimal transport formulation with an optimality prior. Our future work will explore skeleton-based approaches [34, 64] and other applications of our regularized Gromov-Wasserstein optimal transport formulation (e.g., keypoint correspondence [57] and point set registration [60]).
@@ A.1 Hyperparameter Settings
Implementation details are provided in Section 4 of the main paper. Here, we additionally provide the hyperparameter settings used in our experiments across all datasets. Tab. 8 lists the hyperparameters for RGWOT, including the learning rate, optimizer, window size, and other relevant hyperparameters.
@@
Optimizer Adam [39]
@@
A.2 Quantitative Results on All Subtasks of the Egocentric and Third-Person Datasets
Table 9: Results on individual subtasks of EGTEA-GAZE+ [46] dataset.
@@ A.2 Quantitative Results on All Subtasks of the Egocentric and Third-Person Datasets
Table 10: Results on individual subtasks of CMU-MMAC [15] dataset.
@@
Table 11: Results on individual subtasks of ProceL [22] dataset.
@@
Table 12: Results on individual subtasks of CrossTask [74] dataset.
@@
Tabs. 9 and 10 present results on all subtasks of the egocentric datasets CMU-MMAC [15] and EGTEA-GAZE+ [46]. Corresponding results for third-person datasets, including ProceL [22] and CrossTask [74], are provided in Tabs. 11 and 12. Our detailed evaluation spans a wide range of scenarios, offering a detailed assessment of model performance from different viewpoints. These results highlight the robustness and versatility of our approach in addressing diverse video types and tasks, contributing to progress in procedure learning and related fields.

### v2

@@ 5 Conclusion
We present a self-supervised procedure learning framework for recognizing key steps and their order from a collection of unlabeled procedural videos. In particular, our approach leverages a fused Gromov-Wasserstein optimal transport formulation with a structural prior to learn frame-to-frame correspondences between videos while tackling order variations, background/redundant frames, and repeated actions. Furthermore, we empirically examine possible trivial solutions when optimizing solely for the above temporal alignment term and integrate a contrastive regularization term which encourages various frames to be mapped to various points in the embedding space to prohibit the collapse to degenerate solutions. Lastly, extensive evaluations on large-scale egocentric (i.e., EgoProceL) and third-person (i.e., ProCeL and CrossTask) benchmarks are performed to show our superior performance over previous methods, including OPEL which is based on a classical Kantorovich optimal transport formulation with an optimality prior. Our future work will focus on deriving theoretical analyses of degenerate solutions, utilizing alternative regularizations (e.g., TCLR [dave2022tclr]), and explore skeleton-based approaches [hyder2024action, tran2024learning].
@@ Appendix A Supplementary Material
In this supplementary material, we first provide hyperparameter settings of our RGWOT approach across all datasets in Sec. A.1. Next, we present sensitivity analyses of the margin $\lambda$ and weight $\beta$ of the C-IDM regularization, as well as the Gromov-Wasserstein weight $\alpha$, structural prior radius $r$, and temporal prior weight $\rho$ in Sec. A.2, followed by run time comparisons of our RGWOT approach and competing methods including VAOT [ali2025joint] and OPEL [chowdhuryopel] in Sec. A.3. We further provide an empirical justification of degenerate solutions in Sec. A.4. Finally, we present quantitative results on all subtasks of egocentric and third-person datasets in Sec. A.5.
@@ A.1 Hyperparameter Settings
Implementation details have been provided in Sec. 4 of the main paper. Here, we additionally provide hyperparameter settings of our RGWOT approach across all datasets. Tab. 8 lists hyperparameter settings for RGWOT, including the learning rate, optimizer, window size, and other relevant hyperparameters. Note that we use the same hyperparameter settings across all datasets.
@@
Optimizer Adam [kingma2014adam]
@@
A.2 Sensitivity Analyses
(a) Gromov-Wasserstein weight $\alpha$.
(b) Structural prior radius $r$.
(c) Temporal prior weight $\rho$.
(d) Contrastive regularization margin $\lambda$.
Figure 6: Sensitivity analyses of Gromov-Wasserstein weight $\alpha$, structural prior radius $r$, temporal prior weight $\rho$, and contrastive regularization margin $\lambda$.
In addition to the sensitivity analysis of $K$ in Sec. 4.2 of the main paper, we further conduct sensitivity analyses on some key hyperparameters of our RGWOT approach, including the Gromov–Wasserstein weight $\alpha$, structural prior radius $r$, temporal prior weight $\rho$, as well as the margin $\lambda$ and weight $\beta$ of the C-IDM regularization. First of all, Fig. 6 presents sensitivity analyses of the Gromov–Wasserstein weight $\alpha$, structural prior radius $r$, temporal prior weight $\rho$, and margin $\lambda$ on the EgoProceL [bansal2022my] dataset (i.e., PC Assembly). Each subfigure illustrates how the Precision, Recall, F1, and IoU metrics change as the corresponding hyperparameter is adjusted. Overall, RGWOT demonstrates mostly stable performance across the studied ranges, with the performance metrics reaching their peaks at $\alpha=0.3$, $r=0.02$, $\rho=0.35$, and $\lambda=2$.
Next, we study the effects of the C-IDM regularization by varying its weight $\beta$ in Eq. 7 of the main paper and include the results on the ProceL [elhamifar2020self] dataset (i.e., make_smoke_salmon_sandwich) in Tab. 9. It is evident from the results that when $\beta$ is small (e.g., $\beta=0.01$), the learning collapses and RGWOT yields degenerate solutions, where all frames are mapped to a small cluster in the embedding space, and hence an entire video is assigned to a single key steps (see Fig. 3 for examples). Furthermore, when $\beta$ increases (e.g., $\beta=0.05$), the above issue is prevented. Finally, the best performance is achieved when $\beta=1$, indicating balancing weights for the video alignment loss and the contrastive regularization.
$\beta$ Precision Recall F1 IoU
0.01 X X X X
0.05 38.51 40.34 39.40 24.02
0.1 40.49 42.68 41.56 26.73
1 42.61 44.73 43.64 28.22
10 36.41 44.87 40.20 25.33
50 35.41 40.97 37.99 23.54
Table 9: Sensitivity analysis of the contrastive regularization weight $\beta$. Best results are in bold, while second best are underlined. ‘X’ denotes degenerate results.
A.3 Run Time Comparisons
In this section, we evaluate the efficiency (in terms of training times) of our RGWOT approach and competing methods including OPEL [chowdhuryopel] and VAOT [ali2025joint]. In particular, we train all methods on the EgoProceL [bansal2022my] dataset (i.e., PC Disassembly) under identical experimental conditions (e.g., with 10,000 training epochs using two NVIDIA RTX 5090 GPUs). Tab. 10 presents the results. Thanks to having few losses and regularizers, our RGWOT approach is (notably) more efficient than OPEL [chowdhuryopel], i.e., for training, OPEL [chowdhuryopel] needs $206$ minutes vs. $161$ minutes for RGWOT. Thus, our approach is not only more accurate but also more efficient than OPEL [chowdhuryopel]. In addition, RGWOT has similar training time as VAOT [ali2025joint], i.e., for training, VAOT [ali2025joint] requires $164$ minutes vs. $161$ minutes for RGWOT, suggesting that adding contrastive regularization yields little extra computation.
Table 10: Run time comparisons.
Method Training Time
OPEL [chowdhuryopel] 206 (mins)
VAOT [ali2025joint] 161 (mins)
RGWOT (Ours) 164 (mins)
A.4 Degenerate Solution Justification
Figure 7: t-SNE [maaten2008visualizing] plots of the EgoProceL [bansal2022my] dataset. Frame order from first to last is indicated by a gradient from light to dark green.
Figure 8: t-SNE [maaten2008visualizing] plots of the ProceL [elhamifar2020self] dataset. Frame order from first to last is indicated by a gradient from light to dark green.
As discussed in Sec. 3.1.2 of the main paper, the objective in Eq. 5 minimizes visual and temporal differences between corresponding frames in $X$ and $Y$ only, and thus there is no mechanism to prevent the optimization from collapsing. Degenerate solutions appear consistently on the ProceL [elhamifar2020self] dataset but not on the EgoProceL [bansal2022my] and CrossTask [zhukov2019cross] datasets. It is likely because 1) ADAM is a local optimizer, and 2) datasets exhibit different characteristics. To verify the latter, we extract ResNet [he2016deep] features (which are input to the model) from four EgoProceL videos and four ProceL videos and visualize them using t-SNE [maaten2008visualizing] in Figs. 7 and 8 respectively. We observe that EgoProceL features are more spread out, whereas ProceL features are more concentrated and hence easier to collapse. We believe this empirical evidence provides a convincing explanation and consider a theoretical analysis an interesting avenue for our future work.
A.5 Quantitative Results on All Subtasks of Egocentric and Third-Person Datasets
Tabs. 11 and 12 present quantitative results on all subtasks of egocentric datasets, namely EGTEA-GAZE+ [li2018eye] and CMU-MMAC [de2009guide] respectively. Corresponding results for third-person datasets, including ProceL [elhamifar2020self] and CrossTask [zhukov2019cross], are provided in Tabs. 13 and 14 respectively. Our detailed evaluations span a wide range of scenarios, offering detailed assessments of our model performance from different viewpoints. The results highlight the robustness and versatility of our approach in addressing diverse videos and tasks, contributing to progress in procedure learning and related fields.
Table 11: Results on individual subtasks of the EGTEA-GAZE+ [li2018eye] dataset.
@@ A.5 Quantitative Results on All Subtasks of Egocentric and Third-Person Datasets
Table 12: Results on individual subtasks of the CMU-MMAC [de2009guide] dataset.
@@
Table 13: Results on individual subtasks of the ProceL [elhamifar2020self] dataset.
@@
Table 14: Results on individual subtasks of the CrossTask [zhukov2019cross] dataset.